an empty ballon of mass 17.490 grams is filled with 29.64 litres of helium. what force is necessary to keep the balloon from floating away
1
Expert's answer
2020-09-23T11:48:15-0400

"F_{He}=m_{He}\cdot g=\rho_{He}\cdot V_{He}\cdot g=0.18\cdot0.02964\cdot9.81=0.053(N)"


"F_{bal}=m_{bal}\cdot g=0.01749\cdot9.81=0.172(N)"


"F_B=\rho_{air}\cdot V_{air}\cdot g=1.3\cdot0.02964\cdot 9.81=0.378(N)"


"F=F_B-F_{bal}-F_{Hel}=0.378-0.172-0.053=0.153(N)"